OpenStudy (anonymous):

raffle tickets are being sold for a fundraiser the relationship between n number of raffle tickets sold and a amount of money raised is given by the function below A(n)=3n-20 what is the equation of the inverse function N(a)

OpenStudy (anonymous):

@Directrix can you help me please?

Directrix (directrix):

I use x and y. Start with: y = 3x - 20 For the inverse, interchange the x and y and then solve for the new y.

Directrix (directrix):

x = 3y - 20 3y = x + 20 y = (1/3)x + (20/3) Is that N(a) = ?

Directrix (directrix):

What are the options? I don't know what to code the variable x unless it is a. N(a) = (1/3)a + (20/3) --> Maybe this is it. @kimmykahtt